OpenStudy (anonymous):

What is the variance of the following data? If necessary, round your answer to two decimal places. 13, 11, 4, 5, 6, 9, 10, 12, 15, 16

OpenStudy (anonymous):

16.34 17.15 15.29 14.36

OpenStudy (amistre64):

we need the mean to calculate it

OpenStudy (anonymous):

26?

OpenStudy (amistre64):

we dont really need to do it with this data set, we can work it with an equivalently spaced set; lets see what we get when we subtract 10 from every point. (13, 11, 4, 5, 6, 9, 10, 12, 15, 16)-10 3, 1, -6, -5, -4, -1, 0, 2, 5, 6 3, -4,0, 2 = 1 1/10 for a mean add back 10 to get the original mean of the data set; 101/10

OpenStudy (amistre64):

id prolly rather play with the modified set

OpenStudy (amistre64):

subtract the mean from each point and square the results

OpenStudy (anonymous):

b?

OpenStudy (amistre64):

(3, 1, -6, -5, -4, -1, 0, 2, 5, 6) - 1/10 29/10 9/10 -61/10 -51/10 -41/10 -11/10 -1/10 2, 5, 6) - 1/10 b what? i cant tell what it would be till we get to the end

OpenStudy (amistre64):

the wolf says it aint b

OpenStudy (anonymous):

ok..

OpenStudy (amistre64):

2.9 .9 -6.1 -5.1 -4.1 -1.1 - .1 1.9 4.9 5.9 i wonder if we can approximate this and still get a valid result

OpenStudy (amistre64):

9+1+36+25+16 +1+0+4+25+36 ----------------- 10+1+40+50+52 51+52+53 = 156/10 = 15.6

OpenStudy (amistre64):

the wolf says 15.29; so thats alot easier to deal with lol

OpenStudy (anonymous):

thanks alot for the breakdown!!

OpenStudy (amistre64):

i needed the practice ;) good luck